A bicycle path is 7 miles long. A man walks from one end at the rate of 3 mph. At the same time, a friend bicycles from the other end, traveling at 11 mph. In how many minutes will they meet?





Answer by stanbon(75887)

A bicycle path is 7 miles long. A man walks from one end at the rate of 3 mph. At the same time, a friend bicycles from the other end, traveling at 11 mph. In how many minutes will they meet?
----------------------------
Walker's DATA:
rate = 3 mph ; distance = x miles ; time = x/3 hrs
-----
Bicycle's DATA:
rate = 11 mph ; distance = (7-x) miles ; time = (7-x)/11 hrs
-------
Equation:
time = time
x/3 = (7-x)/11
11x = 21-3x
14x = 21
x = 3/2 = 1 1/2 hrs
==========================
Cheers,
Stan H.